# Spaces of subgroups of toral groups

### v1

@@ Abstract.
The author is grateful for comments, discussions and related collaborations with S.Balchin, D.Barnes, T.Barthel, M.Kedziorek, L.Pol, J.Williamson. The work is partially supported by EPSRC Grant EP/W036320/1. The author would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Equivariant homotopy theory in context, where later parts of work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1.
@@ 1.A. Contents
$1\longrightarrow\mathbb{T}\longrightarrow G\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1$
@@
and we show (Corollary [cor:finerr]) that $c$ is an isomorphism up to finite indeterminacy: it is almost surjective, and has finite fibres. There are finitely many different fibres, and these are cohomologically determined. We can also understand how other invariants of subgroups (such as the Weyl groups $W_{G}(K)$) vary with $S$.
@@ 1.B. Motivation
For a compact Lie group $G$, it has been conjectured [6] that the category of rational $G$-spectra is Quillen equivalent to differential graded objects of an abelian category $\mathcal{A}(G)$. The abelian model $\mathcal{A}(G)$ is a category of sheaves over the space $\mathfrak{X}_{G}=\mathrm{Sub}(G)/G$ of conjugacy classes of subgroups of $G$. The basic data associated to a subgroup $K$ is a module over the polynomial ring $H^{*}(BW_{G}^{e}(K))$ with an action of the finite group $W_{G}^{d}(K)$ where $W_{G}(K)=N_{G}(K)/K$ is the Weyl group with identity component $W_{G}^{e}(K)$ and component group $W_{G}^{d}(K)$. As the subgroup varies, $H^{*}(BW_{G}^{e}(K))$ gives a ‘sheaf’ of rings, and $W_{G}^{d}(K)$ gives a ‘component structure’. The basic input data is therefore the space $\mathrm{Sub}(G)/G$ together with the Weyl group $W_{G}(K)$ for each subgroup $K$, and the structure beyond this will be made explicit below.
The relevant structure on $\mathfrak{X}_{G}$ comes from it being the Balmer spectrum of finite rational $G$-spectra. The Zariski topology is determined by the Hausdorff metric topology and the cotoral inclusion order111$(K)\leq_{ct}(H)$ if $K$ is conjugate to a subgroup $K^{\prime}$ normal in $H$ with $H/K^{\prime}$ a torus.
@@ 1.D. Associated work in preparation
This series is part of a more general programme. Future installments will be components with Noetherian Balmer spectra [5] and those with centre-free Balmer spectra [11]. An account of the general nature of the models is in preparation [4], and it is hoped that this will be the basis of the proof that the category of rational $G$-spectra has an algebraic model in general.
@@ 2.A. Crude partition
$1\longrightarrow\mathbb{T}\longrightarrow G\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1$
@@ 2.B. Subgroups of a torus
$\Lambda^{0}=H^{1}(\mathbb{T};\mathbb{Z})=\mathrm{Hom}(H_{1}(\mathbb{T};\mathbb{Z}),\mathbb{Z})\mathrm{Hom}(\Lambda_{0},\mathbb{Z})$
, so that $\Lambda^{0}$ and the toral lattice $\Lambda_{0}$.
@@ 2.C. A little more duality
[…]
@@ 3.A. Splittings
$1\longrightarrow S\longrightarrow H\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1$
@@ 3.B. Subgroups up to conjugacy
Proof : First, we choose a representative splitting $\sigma_{0}$, with $\delta f_{\sigma_{0}}=0$, giving $[f_{\sigma_{0}}]=\epsilon(G)$ in $H^{2}(W;\mathbb{T})$, and since we are assuming $\epsilon(G)$ lifts to $H^{2}(W;S)$, we may assume that $f_{\sigma_{0}}$ takes values in $S$, so that $\sigma_{0}$ gives a subgroup $H=H(S,\sigma_{0})$ and $[f_{\sigma_{0}}]=\epsilon(H)$ in $H^{2}(W;S)$.
Any other splitting $\sigma$ takes the form $\sigma=\sigma_{0}\cdot g$ for $g\in C^{1}(W;\mathbb{T})$. Since $f_{\sigma g}=f_{\sigma}+\delta g$ by Lemma 3.2 (iii), any element of this form satisfies $\delta f_{\sigma}=0$ and $[f_{\sigma}]=\epsilon(G)$.
@@ 3.C. The ramification sequence
[…]
@@
[…]
@@ 3.D. Variation of conjugacy classes
We observe that $H^{1}(W;\mathbb{T}/S)$ is not a continuous function of $S$. Indeed, we may suppose that $G$ is the 2-torus with the group $W$ of order 2 exchanging factors. This means $H_{1}(\mathbb{T})=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ and we take
@@
Now take $S^{n}_{i}=\ker(T^{2}\longrightarrow(\Lambda^{n}_{i})^{*})$. We find $S^{n}_{1}\longrightarrow T^{2}$ (with $H^{1}(W;T^{2}/S^{n}_{1})$ being the sequence constant at $\mathbb{Z}/2$) whilst $S^{n}_{2}\longrightarrow T^{2}$, (with $H^{1}(W;T^{2}/S^{n}_{2})$ being the sequence constant at 00).
@@
This statement explains why we organize our depiction of the space $\mathfrak{X}_{G}$ by using $W$-$\mathrm{Sub}(\Lambda^{0})$.
@@ 4.A. Normalizers direct
(1) $\begin{gathered}H(S,\sigma)\unlhd H(S^{\prime},\sigma)\mbox{ if $S^{\prime}\subseteq S^{+}$ and then }H(S^{\prime},\sigma)/H(S,\sigma)\cong S^{\prime}/S.\end{gathered}$
@@ 4.B. Normalizers by Pontrjagin duality
We will be explicit in the case $\mathbb{T}$ is of rank 2 and where $W$ is cyclic generated by $g$ but the method is perfectly general. First of all, we choose bases so that $(T^{2})^{*}=\mathbb{Z}^{2}$, thus $T^{2}=(\mathbb{Z}^{2})^{*}$ and we will write $\phi:\mathbb{Z}^{2}\longrightarrow T$ for an element with $\phi(1,0)=w,\phi(0,1)=z$. The last piece of notation is that in this basis
$g=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right).$
Now suppose $S$ is finite and $\Lambda^{S}$ has basis $\lambda_{1}=(x_{1},y_{1}),\lambda_{2}=(x_{2},y_{2})$. The condition that $\phi$ lies in $S$ is that $\phi(\lambda_{1})=\phi(\lambda_{2})=1$, which is to say
$w^{x_{1}}z^{y_{1}}=w^{x_{2}}z^{y_{2}}=1.$
@@
[…]
If $w=e^{2\pi ir},z=e^{2\pi is}$ we may rewrite this in a purely additive setting:
$(r\;\;s)\left(\begin{array}[]{cc}x_{1}&x_{2}\\ y_{1}&y_{2}\end{array}\right)\in\mathbb{Z}^{2}.$
Moving on to $\Lambda^{S}_{+}$, the condition for membership is that $\phi(x,y):=\theta g(x,y)/\theta(x,y)$ lies in $S$, which we may check by evaluating on the generators of $\Lambda$: we require
$\theta g(x_{1},y_{1})=\theta(x_{1},y_{1})\mbox{ and }\theta g(x_{2},y_{2})=\theta(x_{2},y_{2}).$
@@
$(r\;\;s)\left(\begin{array}[]{cc}a-1&b\\ c&d-1\end{array}\right)\left(\begin{array}[]{cc}x_{1}&x_{2}\\ y_{1}&y_{2}\end{array}\right)$
is an integer vector. The order of the group $S^{+}/S$ (if finite) is the determinant of the matrix
$M:=\left(\begin{array}[]{cc}a-1&b\\ c&d-1\end{array}\right).$
The structure generally depends on the lattice $\Lambda^{S}$, but if $\det(M)$ is not divisible by the square of any prime it is cyclic of order $|\det(M)|$. In any case $S^{+}$ corresponds to the lattice $\Lambda^{S}_{+}=M\Lambda^{S}$. If $W$ is not cyclic $\Lambda^{S}_{+}$ will be the interesection of the lattices obtained from the generators $g$.
@@
(i) In rank 1, if $W=\langle g\rangle$ is of order 2 and $g$ acts as $-1$ and $S$ is cyclic of order $m$ then $\Lambda^{S}=m\mathbb{Z}$ and the condition for membership of $S$ is
@@
The Weyl group has identity component $T$, but its structure depends on the lattice (see Section 8).
@@
and the Weyl group is $C_{2}\times C_{2}$.
@@
and the Weyl group is of order 3.
@@ 5.B. Tori
[…]
@@
This is sufficient to give an estimate for arbitrary subgroups.
We may now prove the proposition.
@@
Since the number of vectors with all terms $<N$ is finite for any $N$, this base of neighbourhoods generates the same topology as the cofinite base of neighbourhoods so the topology is entirely determined by containment of subgroups.
@@ 6.A. The method
[…]
@@ 6.B. Catalogue of examples
§$\begin{array}[]{|lcc|c|cc|c|cc|}\S&W&\Lambda_{0}&\Lambda_{S}&H^{2}(W;\Lambda_{0})&H^{2}(W;\Lambda_{S})&H^{2}(W;S)&H^{3}(W;\Lambda_{0})&H^{3}(W;\Lambda_{S})\\ \hline\cr&&&&A&B&C&D&E\\ \hline\cr 8&C_{2}&\mathbb{Z}\oplus\tilde{\mathbb{Z}}&\mathbb{Z}\oplus\tilde{\mathbb{Z}}&2&2&&2&2\\ &&&\mathbb{Z}W&2&0&*&2&0\\ \hline\cr 9&C_{2}&\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}&\tilde{\mathbb{Z}}\oplus\tilde{\mathbb{Z}}&0&0&&2^{2}&2^{2}\\ \hline\cr 10&C_{2}&\mathbb{Z}W&\mathbb{Z}\oplus\tilde{\mathbb{Z}}&0&2&*&0&2\\ &&&\mathbb{Z}W&0&0&*&0&0\\ \hline\cr 11&C_{2}\times C_{2}&A_{1}\times A_{1}&\tilde{\mathbb{Z}}_{1}\oplus\tilde{\mathbb{Z}}_{2}&2^{2}&2^{2}&&2^{4}&2^{4}\\ &&&FCC_{A_{1}\times A_{1}}&2^{2}&0&&2^{4}&2\\ \hline\cr 12&C_{2}\times C_{2}&\mathrm{Cub}_{\delta}&\mathrm{Cub}_{\delta}&0&2^{2}&&2&2^{4}\\ &&&FCC_{\delta}&0&0&&2&2\\ \hline\cr 13&C_{4}&\mathrm{Cub}_{\pi/2}&\mathrm{Cub}_{\pi/2}&0&0&&2&2\\ &&&FCC_{\pi/2}&0&0&&2&2\\ \hline\cr 14&D_{8}&B_{2}&B_{2}&2&2&&2&2\\ &&&FCC_{B_{2}}&2&2&&2&2\\ \hline\cr 15&C_{3}&\mathbb{Z}[\omega]&\mathbb{Z}[\omega]&0&0&&3&3\\ \hline\cr 16&D_{6}&A_{2}&\mathbb{Z}[\omega]^{\prime}&0&0&*&0&0\\ \hline\cr&&\mathbb{Z}[\omega]^{\prime\prime}&\mathbb{Z}[\omega]^{\prime\prime}&0&0&&3&3\\ \hline\cr 17&D_{12}&G_{2}&\mathbb{Z}[\omega]^{\prime}&0&0&*&0&0\\ \hline\cr\end{array}$

### v2

@@ Abstract.
The author is grateful for comments, discussions and related collaborations with S.Balchin, D.Barnes, T.Barthel, Z.Huang, M.Kedziorek, L.Pol, J.Williamson. The work is partially supported by EPSRC Grant EP/W036320/1. The author would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programmes Topology, representation theory and higher structures and Equivariant homotopy theory in context, where later parts of work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1. The author declares there are no competing interests.
@@ 1.A. Contents
$1\longrightarrow\mathbb{T}\longrightarrow G\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1.$
@@
and we show (Corollary 3.5) that $c$ is an isomorphism up to finite indeterminacy: it is almost surjective, and has finite fibres. There are finitely many different fibres, and these are cohomologically determined. We can also understand how other invariants of subgroups (such as the Weyl groups $W_{G}(K)$) vary with $S$.
@@ 1.B. Motivation
For a compact Lie group $G$, it has been conjectured [6] that the category of rational $G$-spectra is Quillen equivalent to differential graded objects of an abelian category $\mathcal{A}(G)$. The abelian model $\mathcal{A}(G)$ is a category of sheaves over the space $\mathfrak{X}_{G}=\mathrm{Sub}(G)/G$ of conjugacy classes of subgroups of $G$. The basic data associated to a subgroup $K$ is a module over the polynomial ring $H^{*}(BW_{G}^{e}(K))$ with an action of the finite group $W_{G}^{d}(K)$ where $W_{G}(K)=N_{G}(K)/K$ is the Weyl group with identity component $W_{G}^{e}(K)$ and component group $W_{G}^{d}(K)$. As the subgroup varies, $H^{*}(BW_{G}^{e}(K))$ gives a ‘sheaf’ of rings, and $W_{G}^{d}(K)$ gives a ‘component structure’. The basic input data is therefore the space $\mathrm{Sub}(G)/G$ together with the Weyl group $W_{G}(K)$ for each subgroup $K$.
[…]
@@ 1.D. Associated work in preparation
This series is part of a more general programme. Future installments will be blocks with Noetherian Balmer spectra [5] and blocks with all Weyl groups finite [11]. The implications of the methods here for the structure of general blocks are explained in [13]. Building on this, an account of the general nature of the models is in preparation [4], and it is hoped that this will be the basis of the proof that the category of rational $G$-spectra has an algebraic model in general.
@@ 2.A. Crude partition
$1\longrightarrow\mathbb{T}\longrightarrow G\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1.$
@@ 2.B. Subgroups of a torus
$\Lambda^{0}=H^{1}(\mathbb{T};\mathbb{Z})=\mathrm{Hom}(H_{1}(\mathbb{T};\mathbb{Z}),\mathbb{Z})=\mathrm{Hom}(\Lambda_{0},\mathbb{Z}),$
so that the lattices $\Lambda^{0}$ and $\Lambda_{0}$ are dual.
@@ 2.C. A little more duality
${H^{2}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{3}(W;\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{2}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{3}(W;\Lambda_{S})}$
@@ 3.A. Splittings
$1\longrightarrow S\longrightarrow H\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}W\longrightarrow 1.$
@@ 3.B. Subgroups up to conjugacy
Proof : First, we choose a representative splitting $\sigma_{0}$, with $\delta f_{\sigma_{0}}=0$, giving $f_{\sigma_{0}}]=\epsilon(G)$ in $H^{2}(W;\mathbb{T})$, and since we are assuming $\epsilon(G)$ lifts to $H^{2}(W;S)$, we may assume that $f_{\sigma_{0}}$ takes values in $S$, so that $\sigma_{0}$ gives a subgroup $H=H(S,\sigma_{0})$ and $f_{\sigma_{0}}]=\epsilon(H)$ in $H^{2}(W;S)$.
Any other splitting $\sigma$ takes the form $\sigma=\sigma_{0}\cdot g$ for $g\in C^{1}(W;\mathbb{T})$. Since $f_{\sigma g}=f_{\sigma}+\delta g$ by Lemma 3.2 (iii), any element of this form satisfies $\delta f_{\sigma}=0$ and $f_{\sigma}]=\epsilon(G)$.
@@ 3.C. The ramification sequence
$\mathrm{Split}(G|S)\longrightarrow\mathrm{Sub}(G|S)\longrightarrow\mathrm{Sub}(G|S)/G\longrightarrow\mathrm{Sub}(G|S)/\mbox{ext-iso}$
@@
${\mathrm{Split}(G|S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Split}(G|S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Sub}(G|S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Split}(G|S)/C^{1}(W;S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${Z^{1}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\kappa_{0}}$${\mathrm{Sub}(G|S)/G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Split}(G|S)/C^{1}(W;S),B^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{1}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\kappa_{0}}$${\mathrm{Sub}(G|S)/\mbox{ext-iso}}$${\mathrm{Split}(G|S)/C^{1}(W;S),Z^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{1}(W;\mathbb{T}/S)/\mathrm{im}H^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\kappa_{0}}$
@@ 3.D. Variation of conjugacy classes
We observe that $H^{1}(W;\mathbb{T}/S)$ is not a continuous function of $S$. Indeed, we may suppose that $G$ is the 2-torus with the group $W$ of order 2 acting trivially on one factor and as inversion on the other. This means $H_{1}(\mathbb{T})=\mathbb{Z}\oplus\tilde{\mathbb{Z}}$ and we take
@@
Now take $S^{n}_{i}=\ker(T^{2}\longrightarrow(\Lambda^{n}_{i})^{*})$. We find $S^{n}_{1}\longrightarrow T^{2}$ (with $H^{1}(W;T^{2}/S^{n}_{1})$ being the sequence constant at $\mathbb{Z}/2$) whilst $S^{n}_{2}\longrightarrow T^{2}$, (with $H^{1}(W;T^{2}/S^{n}_{2})$ being the sequence constant at 0).
@@
Based on this, we depict the space $\mathfrak{X}_{G}$ lying over $W$-$\mathrm{Sub}(\Lambda^{0})$.
@@ 4.A. Normalizers direct
[…]
@@ 4.B. Normalizers by Pontrjagin duality
We suppose $\mathbb{T}$ is a torus of rank $r$ and start with the case that $W$ is cyclic generated by $g$. First of all, we choose bases so that $\mathbb{T}^{*}=\mathbb{Z}^{r}$, and we identify $\mathbb{T}$ with its double dual $\mathbb{T}\cong(\mathbb{Z}^{r})^{*}$. We will write $\phi:\mathbb{Z}^{r}\longrightarrow T$ for an element of $S$ and name the images of the standard basis $\phi(e_{i})=z_{i}\in T$ for $1\leq i\leq r$. The last piece of notation is that in this basis $g$ is given by the matrix $A(g)$.
Now suppose $S$ is of dimension rkr-k and $\Lambda^{S}$ has basis $\lambda^{1},\ldots,\lambda^{k}$, and we express the $\lambda^{i}$ in terms of the standard basis: $\lambda^{i}=(x^{i}_{1},\ldots,x^{i}_{r})$. We write $L$ for the matrix with columns $\lambda^{1},\ldots,\lambda^{k}$. The condition that $\phi$ lies in $S$ is that $\phi(\lambda^{i})=1$ for $1\leq i\leq k$, which is to say
$w_{1}^{x^{i}_{1}}w_{2}^{x^{i}_{2}}\cdots w_{r}^{x^{i}_{r}}=1\mbox{ for }i=1,\ldots,k.$
@@
$(w_{1},\ldots,w_{r})L=(1,\ldots,1).$
Converting to additive notation, if $w_{j}=e^{2\pi is_{j}}$ for $1\leq j\leq r$, we may rewrite this in a purely additive setting:
$(s_{1},\ldots,s_{r})L\in\mathbb{Z}^{2}.$
Moving on to $S_{+}$, the condition that $\theta$ lies in $S_{+}$ is that $\phi:=\theta^{-1}\theta^{g}$ lies in $S$. Including variables for the functions on $\mathbb{Z}^{r}$ this states that $\phi(x_{1},\ldots,x_{r})=\theta A(g)(x_{1},\ldots,x_{r})/\theta(x_{1},\ldots,x_{r})$ lies in $S$, which we may check by evaluating on the generators of $\Lambda^{S}$: we require
$\theta A(g)\lambda^{i}=\theta\lambda^{i}\mbox{ in $T$ for }1\leq i\leq k$
@@
$(s_{1},\ldots,s_{r})M(g)L$
is an integer vector where M(g)=A(g)IM(g)=A(g)-I. In other words $S_{+}$ is dual to the lattice $M(g)L$. If $S$ is finite (so that $L$ is a non-singular square matrix) the order of $S^{+}/S$ (if finite) is $|\det(M(g))|$, but the structure will depend on the lattice and the actual matrix.
If $W$ is not cyclic, we require that the condition holds each generator $g$ of $W$, so we find $S_{+}$ is dual to the lattice $\Lambda^{S}_{+}=\sum_{g}M(g)L$.
@@
(i) In rank 1, if $W=\langle g\rangle$ is of order 2 and $g$ acts as 1-1 and $S$ is cyclic of order $m$ then $\Lambda^{S}=m\mathbb{Z}$ and the condition for membership of $S$ is
@@
The Weyl group of each finite $S$ has identity component $T$, but its structure depends on the lattice (see Section 8).
@@
and the Weyl group of any finite $S$ is $C_{2}\times C_{2}$.
@@
and the Weyl group of every finite $S$ is of order 3.
@@ 5.B. Tori
In fact we only need to estimate the distance between a subgroup $S_{*}$ and a codimension 1 subgroup $S$ of $\mathbb{T}$. One should think of the a circle of slope $p/q$ (in lowest terms) in a 2-torus will be within about $1/2q$ of every other point. For a codimension 1 subgroup, $\Lambda^{S}=\langle\theta\rangle$ where $\theta=(a_{1},\ldots,a_{r})\neq 0$, we have $\mu(S)=\max\{|a_{i}|\;|\;0\leq i\leq r\}.$
@@
This is sufficient to give an estimate for arbitrary subgroups, and we may now prove the proposition.
@@
Since the number of vectors with all terms <N<N is finite for any $N$, this base of neighbourhoods generates the same topology as the cofinite base of neighbourhoods so the topology is entirely determined by containment of subgroups.
@@ 6.A. The method
    ${H^{1}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a_{S}^{1}}\scriptstyle{\cong}$${H^{1}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{2}(W;S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{2}(W;\mathbb{T})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{a_{S}^{2}}\scriptstyle{\cong}$${H^{2}(W;\mathbb{T}/S)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${H^{2}(W;\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k^{S}_{*}}$${H^{2}(W;\Lambda_{S})}$${H^{3}(W;\Lambda_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k^{S}_{*}}$${H^{3}(W;\Lambda_{S})}$
@@ 6.B. Catalogue of examples
[…]